# End-to-End Demonstration for CubeSatellite Quantum Key Distribution

### v1

@@ 1 Introduction
Quantum technology has already started a technological revolution in the fields of communication [1], sensing [2, 3] and computing due to its enhanced capabilities. Quantum networks are moving step by step from the laboratory to end-user applications due to the threat of emerging quantum computers to existing cryptographic systems and the quest for unconditional security. A main difficulty to establish a global scale quantum network is exponential loss in optical fibres which limits the direct communication distance to less than 1000 km [4, 5, 6, 7]. Quantum repeaters are a good option for extending communication distances, but global extension through a large number of repeaters is not realistic [8]. In addition, establishing fiber-based networks in areas isolated from the fibre backbone is impractical due to geographical factors and the absence of infrastructure, while mobile platforms require more flexibility of communication locations.
@@
The near vacuum and microgravity environment of space makes it a potential way to challenge fundamental physics, including quantum technology [9]. Early feasibility studies showed that satellite-based quantum communications can transmit photons over distances of several thousand kilometres [10, 11], making it ideal to create worldwide key distribution by linking into fibre QKD networks for local distribution [12]. The idea was discussed for several years and the interest reached a peak with the successful launch and demonstration of the Micius satellite [13]. This first demonstration was a large and complex satellite weighing some 500kg requiring a dedicated, thus expensive, launch. As technology advances, there is an expectation that in-orbit demonstrations of quantum key distribution using nanosatellites will be possible due to their compact design, lower cost and the feasibility of establishing constellations [14].
@@
However, the limited space and scarce power budget raise a series of challenges for the payload implementation. A Cube-Satellite’s size is from 1 U to 16 U1U is a standard dimension (Units or “U”) of 10 cm x 10 cm x 10 cm, which requires the payload to be compact. It is a particular challenge for the optical component miniaturisation, electronic integration and electromagnetic compatibility design. The solar panel size and the corresponding generated electric power decreases with a smaller satellite, which requires a more elegant power distribution and higher efficiency device selection. This limited resource exacerbates the difficulty of acquisition, pointing and tracking (APT), as well as the T&S.
@@
In this work, we built up an end-to-end testbench used for demonstrating satellite-based QKD performance in the laboratory. The system used weak coherent pulse as quantum source. The T&S between the satellite and optical ground station (OGS) is achieved using a hybrid de Bruijn code (HDBC) implemented by a high power pulsed laser diode  [15]. In the free space test channel, we use a variable ND filter to simulate the channel loss changing when the satellite is in orbit. We perform a series experiments to simulate the scenario in which the channel loss, and background noise is changing. Also we investigate the effect on secure key generation when selecting different gate widths in the postprocessing. Finally, the finite key length and loss budget is discussed with suggested improvements to the future design for satellite flight model development. The transmitter developed here is planned to be implemented in three future nano-satellite missions: in a UK-Canada add-on to the QEYSSAT mission  [16], in the ROKS mission  [17] and in the UK Quantum Communication Technology Hub Mission  [18], the design work will mainly focus on the Hub Mission.
@@ 2 Challenge and System Design
Link Efficiency. Compared with the maximum communication distance of 1000 km of optical fiber [7], satellite QKD can realize communication ranges of several thousand kilometers. However, due to the beam diffraction limit and limited receiver aperture, both beacon and quantum beam suffer diffraction loss. In the Hub Mission for example, geometric losses can rise to 27.7 dB at the low elevation angles (10 degrees) and 500 km altitude when quantum beam divergence is specified as 10 microrad ($1/e^{2}$)  [18]. Here the downlink T&S beacon is designed to be slightly larger divergence at 18.7 microrad ($1/e^{2}$) to reliably perform the timing and synchronisation, resulting in 5.4 dB increase in geometric losses compared to the quantum signal (ie up to 33.1 dB loss). For system efficiency, as the transmitter can be treated as a trusted node, it is only necessary to ensure that the mean photon number $\mu$ at the telescope’s exit aperture meets the security conditions, which alleviates internal optical losses. However for the beacon, the transmitter internal loss is about 3 dB due to the alignment error and the transmission of the dichroic mirror for combining the beacon and quantum beam. Carbon dioxide, oxygen, nitrogen, water and aerosol/particulate species in Earth atmosphere determine a large variety of absorption and scattering processes. The transmission at 785 nm is 0.56 (2.5 dB loss) and at 905 nm is 0.96 (0.2 dB loss) in zenith. [19, 20] and both are about 7.9 dB loss at a distance of 1700 km, corresponding to a 10 degrees elevation angle. The downlink beam configuration suffers comparatively little loss due to turbulence compared with the uplink beam configuration. The geometric losses depend on the OGS receiver telescope aperture area, and we assume in the Hub mission estimates that a telescope diameter of 70 cm is used. We assume for pointing error losses (including turbulence) will contribute 3 dB loss and OGS internal filtering and coupling will contribute another 5 dB losses, and the detector efficiency of 60% adds another 2.2 dB of loss. So, the total quantum link loss is between 29.8 dB - 45.8 dB while the beacon link efficiency is between 31.7 dB - 50 dB. The loss budget is sumarised in Table. 2 and the loss modelling is presented in the appendix.
Table 1: Loss budget comparison between two missions
ROKs & UK-Canada Mission Quantum (785 nm) Beacon (905 nm)
Tx Telescope Diameter 90 mm 90 mm
OGS Telescope Diameter 432 mm 432 mm
Divergence 10 urad 100 urad
Losses internal to transmitter - 13.2 dBIn order to get enough peak power for reliable SNR in the receiver APD, the beacon source use a surface mount multi-mode laser diode which could provide up to 100W peak power but could only get 0.28 mrad with a 5mm beam size. This will result in a larger divergence compared with the target 100urad. On the other hand, the limited internal aperture (3.6mm) before the transmitted telescope clip part of the beam, which introduce an extra loss.
@@
Losses in fading channel 21.7 dB - 32.3 dBThe minimum and maximum values of the loss correspond to the position of the satellite at 90° elevation angle and the position of the satellite at 10° elevation angle, respectively. 41.7 dB - 52.3 dB66footnotemark: 6
Turbulence and pointing losses 3 dB 3 dB
Atmosphere Absorption 2.5 dB - 7.9 dB66footnotemark: 6 0.17dB - 7.9dB66footnotemark: 6
Losses in OGS 5 dB 5 dB
Quantum Detection efficiency 2.2 dB -
Total Loss 34.4 dB - 50.4 dB66footnotemark: 6 61.1 dB - 79.4 dB66footnotemark: 6
Hubs Mission Quantum (785 nm) Beacon (915 nm)
Tx Telescope Diameter 80 mm 80 mm
OGS Telescope Diameter 700 mm 700 mm
Divergence 10 urad 18.7 urad
Losses internal to transmitter - 3 dB
Losses in fading channel 17.1 dB - 27.7 dB66footnotemark: 6 22.5 dB - 33.1 dB66footnotemark: 6
Turbulence and pointing losses 3 dB 3 dB
Atmosphere Absorption 2.5 dB - 7.9 dB66footnotemark: 6 0.17 dB - 7.9 dB66footnotemark: 6
Losses in OGS 5 dB 3 dB
Quantum Detection efficiency 2.2 dB -
Total Loss 29.8 dB - 45.8 dB66footnotemark: 6 31.7 dB - 50.0 dB66footnotemark: 6
@@
Relative Motion. Currently in-orbit and under-development QKD satellites are all running in low earth orbit (LEO). Unlike geosynchronous earth orbit (GEO) satellites, LEO orbit periods are much faster than the Earth’s rotation period. For a LEO at 500km altitude, its orbital period is about 1.2 hours. This results in continuous relative motion between the satellite and the OGS, which leads, on the one hand to a constant variation in link loss and on the other hand to high demands on the tracking and pointing systems between the satellites and OGS. In addition, due to the varying attitude changes of the satellites relative to the OGS, polarisation compensation is required to correct the polarisation reference between both sides, which is particularly important in the case of the BB84 protocol based on polarisation coding. Last but not least, varying linear relative motion causes the signal to carry the Doppler shift, which not only increases signal jitter but also rises a wavelength shift for both beams. Assuming $\pm$10 km/sec maximum relative velocity this would imply ±$\pm 3\times 10^{-5}$ relative frequency shift [21]. A very narrow bandpass filter might cause transmission variability due to the wavelength shifting from the Doppler Effect, but this is negligible for a filter whose window is $>1$ nm. The effect of Doppler shift on temporal synchronisation using clock recovered from a 100 kHz beacon is also negligible compared with the 100 ps synchronisation jitter requirement[15, 22].
@@
Timing and Sync. To ensure consistent timing between the transmitter and receiver, an elegant time synchronisation mechanism needs to be designed as large and costly devices such as high precision atomic clocks are not suitable for this scenario. Several different approaches have been utilized for T&S over long free space channels, including using pulsed lasers with various modulation patterns, and even combined with global positioning system (GPS) pulse per second (PPS) as a hybrid scheme. However, the $>$ 50 dB loss budget greatly increases the difficulty of precision beacon timing detection. A ultra-weak optical signal detection requires a detector with very high sensitivity and low noise level. A higher detecting bandwidth also provides a better timing jitter due to the sharper rise time. However, due to the limitations of the detector gain–bandwidth product (GBP), sensitivity and bandwidth need to be traded off to achieve the best SNR. Additionally, a large variation of incident power can result in pulse detection failure due to signal saturation. More importantly, the amplitude variation of the converted electronic pulse causes large discrimination errors of up to nanosecond level when using constant threshold signal discrimination, which significant increase the timing jitter and degrades the SNR [23]. Our solution is detailed below and in previous papers [15, 22].
@@ 2.1 Communication Link Design
The BB84 protocol encodes photons using four independent polarisation states, but the laser beam generated by an actual source can never reach 100% linear polarization. When photons coded as H-polarisation are incident on an ideal polarising beam splitter (PBS), a small fraction will inevitably enter into the V polarisation channel and thus contribute to a raw key error rate, denoted as $QBER_{SP}$. Similarly, the imperfection of practical PBS’s introduce unavoidable QBER as well denoted as $QBER_{PBS}$. In practice, The $QBER_{PBS}$ can be estimated based on the extinction ratio of the PBS. The $QBER_{SP}$ is decided by the laser and polarizer. A polarizer with higher extinction ratio gives a lower QBER. These two parts are only determined by the device performance and working environment such as temperature and humidity, rather than the system design parameters, so they are defined as intrinsic QBER, $QBER_{I}$.
$QBER_{I}=F(QBER_{SP},QBER_{PBS})$ (1)
$F$ is the QBER combination function, which is defined as below:
$F(Q_{A},Q_{B})=Q_{A}+Q_{B}-2Q_{A}Q_{B}$ (2)
@@
Where the $QBER_{A}$ is the part resulted from atmosphere, the $QBER_{N}$ is the part resulted from background noise and the $QBER_{DCR}$ is the noise comes from DCR. So, the whole $QBER$ is defined as
$QBER=F(QBER_{I},QBER_{E})$ (4)
@@
In QKD security analysis, any $QBER$ must be considered as resulting from the activity of an eavesdropper, which requires post-processing to perform privacy amplification of the key based on the $QBER$ to ensure adequate security. Therefore, efforts to reduce the $QBER$ can increase the key generation rate, and the optimization process of the system focuses on the $QBER_{E}$ since the $QBER_{I}$ is only determined by the performance of the source, the detector and the optics [30]. Bandpass filters are usually used to suppress background noise, which can greatly reduce all other noise beyond the band of the system signal. Generally speaking, the narrower filter window and the steeper cut-off edge, the better signal to noise ratio (SNR) could be achieved. However, considering the wavelength shift caused by the Doppler effect and temperature, the selection of the bandpass filter needs to be balanced between signal transmission and noise rejection to maximise the overall SKR. Besides, the background noise can be mitigated by gating or time-stamping the detection events. The minimum time interval for gating or time-stamping is set by the timing jitter of the detector, beacon and time-to-digital converter.
@@
As optical wavelengths are several orders of magnitude shorter than radio frequency signals, the diffraction limited divergence angles of 10 cm diameter beams are on the order of microradians. This brings optical communication the advantage of higher transmission efficiency compared to microwave communication, but it also leads to the need for an accurate pointing and tracking system to match the fields of view of satellites and OGS [31]. In addition, due to the relative motion of satellites with respect to ground station and the complex and changeable atmospheric environment, traditional clock synchronisation schemes for QKD in free space will lead to large jitter. The lack of synchronised high precision clocks results in inconsistencies between the transmitted and the received sequence and poor discrimination of background counts as a wide time gate is needed. Also, due to the continuous change of the satellite attitude relative to the OGS, providing an identical and stable polarisation reference is also essential to successful communication. Based on these considerations we have developed an elegant solution using a classical bright pulsed laser as beacon [15], and here is the first time it has be used for QKD demonstration. The pulses are designed to be highly synchronised (less than 15 ps) with the quantum signal to indicate the exact time of quantum signal arriving. Each two adjacent beacon pulses represent a bit and the HDBC is modulated onto the pulses train to provide a unique sequence index for each arriving pulse. Finally, the beacon source adopts linearly polarized light with a sufficiently high extinction ratio and is consistent with the polarization state of the H channel, so as to provide a reference for the OGS. Polarisation matching through satellite attitude control can solve the above problems, but it requires accurate, stable and real-time closed-loop feedback control, which dramatically increases the complexity of the system and increases the fuel consumption of the satellite. Additionally it cannot compensate for the error caused by any atmospheric effects. The most widely used method is to polarisation compensate the received signal with a set of automatically rotating waveplates. Using the linear polarisation properties of the beacon light, an algorithm will control the automatic rotation of two quarter-wave plates and one half-wave plate until the intensity of the received light reaches its maximumIt is worth noting that because the beacon light and the signal light use different wavelengths, any birefringence caused by atmospheric effects is not independent of the wavelength, so it may lead to greater compensation errors in the signal at the receiving end..
3 System and Result

### v2

@@ 1 Introduction
Quantum technology has already started a technological revolution in the fields of communication [1], sensing [2, 3] and computing due to its enhanced capabilities. Quantum networks are moving step by step from the laboratory to end-user applications due to the threat of emerging quantum computers to existing cryptographic systems and the quest for unconditional security. A main difficulty to establish a global scale quantum network is exponential loss in optical fibres which limits the direct communication distance to around 1000 km [4, 5, 6, 7, 39]. Quantum repeaters are a good option for extending communication distances, but global extension through a large number of repeaters is not realistic [8]. In addition, establishing fiber-based networks in areas isolated from the fibre backbone is impractical due to geographical factors and the absence of infrastructure, while mobile platforms require more flexibility of communication locations.
@@
The near vacuum and microgravity environment of space makes it a potential way to test fundamental physics, including quantum technology [9]. Early feasibility studies showed that satellite-based quantum communications can transmit photons over distances of several thousand kilometres [10, 11], making it ideal to create worldwide key distribution by linking into fibre QKD networks for local distribution [12]. The idea was discussed for several years and the interest reached a peak with the successful launch and demonstration of the Micius satellite [13]. This first demonstration was a large and complex satellite weighing some 500kg requiring a dedicated, thus expensive, launch. As technology advances, there is an expectation that in-orbit demonstrations of quantum key distribution using nanosatellites will be possible due to their compact design, lower cost and the feasibility of establishing constellations [14].
@@
However, the limited space and scarce power budget raise a series of challenges for the payload implementation. A Cube-Satellite’s size is from 1 U to 16 U1111U is a standard dimension (Units or “U”) of 10 cm x 10 cm x 10 cm, which requires the payload to be compact. It is a particular challenge for the optical component miniaturisation, electronic integration and electromagnetic compatibility design. The solar panel size and the corresponding generated electric power decreases with a smaller satellite, which requires a more elegant power distribution and higher efficiency device selection. This limited resource exacerbates the difficulty of acquisition, pointing and tracking (APT), as well as the T&S.
@@
In this work, we built up an end-to-end testbench used for demonstrating satellite-based QKD performance in the laboratory. The system used weak coherent pulse as quantum source. The T&S between the satellite and optical ground station (OGS) is achieved using a hybrid de Bruijn code (HDBC) implemented by a high power pulsed laser diode  [15]. In the free space test channel, we use a variable ND filter to simulate the channel loss changing when the satellite is in orbit. We perform a series experiments to simulate the scenario in which the channel loss, and background noise is changing. Also we investigate the effect on secure key generation when selecting different gate widths in the postprocessing and the finite key length that can be generated. The work described here is the proof of principle experiment which reproduces the losses, and background noise expected in a space channel as modelled in a separate paper [19] using a robust synchronisation scheme  [15] that enables operation in high loss and automatically corrects the varying Doppler shifts between the clocks. We are able to optimise gate widths and photon numbers to exchange keys over high loss links using electronics and hardware which we are now developing to fit a $\frac{1}{2}$ system [34]. The transmitter and post processing software developed here is planned to be implemented in the UK Quantum communications hub SPOQC Mission [16] where collaborators are developing a 12U nano-satellite with telescope [40] and full acquisition, pointing and tracking. The success of this work has also led to further planned missions including the UK-Canada add-on QEYSSAT mission [17] and the ROKS mission  [18].
@@ 2 Challenge and System Design
Link Efficiency. Compared with the maximum communication distance around 1000 km of optical fiber [7], satellite QKD can realize communication ranges of several thousand kilometers. However, due to the beam diffraction limit and limited receiver aperture, both beacon and quantum beam suffer diffraction loss. In the SPOQC Mission for example, geometric losses can rise to 27.7 dB at the low elevation angles (10 degrees) and 500 km altitude when quantum beam divergence is specified as 10 microrad ($1/e^{2}$)  [19]. Here the downlink T&S beacon is designed to be slightly larger divergence at 18.7 microrad ($1/e^{2}$) to reliably perform the timing and synchronisation, resulting in 5.4 dB increase in geometric losses compared to the quantum signal (ie up to 33.1 dB loss). For system efficiency, as the transmitter can be treated as a trusted node, it is only necessary to ensure that the mean photon number $\mu$ at the telescope’s exit aperture meets the security conditions, which alleviates internal optical losses. However for the beacon, the transmitter internal loss is about 3 dB due to the alignment error and the transmission of the dichroic mirror for combining the beacon and quantum beam. Carbon dioxide, oxygen, nitrogen, water and aerosol/particulate species in Earth atmosphere determine a large variety of absorption and scattering processes. The transmission at 785 nm is 0.56 (2.5 dB loss) and at 905 nm is 0.96 (0.2 dB loss) in zenith. [20, 21] and both are about 7.9 dB loss at a distance of 1700 km, corresponding to a 10 degrees elevation angle. The downlink beam configuration suffers comparatively little loss due to turbulence compared with the uplink beam configuration. The geometric losses depend on the OGS receiver telescope aperture area, and we assume in the Hub mission estimates that a telescope diameter of 70 cm is used. In the downlink scheme, only the end of the optical path (close to the surface of the Earth) is affected by the turbulence, so the pointing error resulting from turbulence will lead to some additional loss. Here we assume pointing error losses (including from turbulence) will contribute 3 dB loss and OGS internal filtering and coupling will contribute another 3.8 dB losses, and the detector efficiency of 60% adds another 2.2 dB of loss. So, the total quantum link loss is between 28.6 dB and 44.6 dB while the beacon link loss is between 31.7 dB and 50 dB. The loss budget is summarised in Table. 1 and the loss modelling is presented in the appendix.
@@
Table 1: Loss budget comparison between two missions  [^3] The minimum and maximum values of the loss correspond to the satellite position at 90 deg elevation angle and 10 deg elevation angle with 500 km orbit, respectively.
ROKs Q (785 nm) SPOQC Q(785 nm) ROKs B (905 nm) SPOQC B (915 nm)
Tx Telescope Diameter (exp) 90 mm 80 mm 90mm 80 mm
OGS Telescope Diameter (exp) 432 mm 700 mm 432 mm 700 mm
Divergence (sim) 10 urad 10 urad 100 urad 18.7 urad
Losses internal to transmitter (exp) 0 dB 0 dB 13.2 dB [^4] In order to get enough peak power for reliable SNR in the receiver APD, the beacon source uses a surface mount multi-mode laser diode which could provide up to 100W peak power but could only get 0.28 mrad with a 5mm beam size. This will result in a larger divergence compared with the target 100urad. On the other hand, the limited internal aperture (3.6mm) before the transmitted telescope clips part of the beam, which introduce an extra loss. 3 dB
Geometric Losses in optical channel (sim) 21.7 dB - 32.3 dB 17.1 dB - 27.7 dB 41.7 dB - 52.3 dB 22.5 dB - 33.1 dB
Turbulence and pointing losses [13] 3 dB 3 dB 3 dB 3 dB
Atmosphere Absorption (ref) 2.5 dB - 7.9 dB 2.5 dB - 7.9 dB 0.17dB - 7.9dB 0.17 dB - 7.9 dB
Losses in OGS (exp) 3.8 dB 3.8 dB 3 dB 3 dB
Quantum Detection efficiency (exp) 2.2 dB 2.2 dB - -
Total Loss 33.2 dB - 49.2 dB 28.6 dB - 44.6 dB 61.1 dB - 79.4 dB 31.7 dB - 50.0 dB
@@
Relative Motion. Currently in-orbit and under-development QKD satellites are all running in low earth orbit (LEO). Unlike geosynchronous earth orbit (GEO) satellites, LEO orbit periods are much faster than the Earth’s rotation period. For a LEO at 500 km altitude, its orbital period is about 1.2 hours. This results in continuous relative motion between the satellite and the OGS, which leads, on the one hand to a constant variation in link loss and on the other hand to high demands on the tracking and pointing systems between the satellites and OGS. In addition, due to the varying attitude changes of the satellites relative to the OGS, polarisation compensation is required to correct the polarisation reference between both sides, which is particularly important in the case of the BB84 protocol based on polarisation coding. Last but not least, varying linear relative motion causes the signal to carry the Doppler shift, which not only increases signal jitter but also rises a wavelength shift for both beams. Assuming $\pm$10 km/sec maximum relative velocity this would imply ±$\pm 3\times 10^{-5}$ relative frequency shift [22]. A very narrow bandpass filter might cause transmission variability due to the wavelength shifting from the Doppler Effect, but this is negligible for a filter whose window is $>1$ nm. The effect of Doppler shift on temporal synchronisation using clock recovered from a 100 kHz beacon is also negligible compared with the 100 ps synchronisation jitter requirement.
@@
Timing and Sync. To ensure consistent timing between the transmitter and receiver, an elegant time synchronisation mechanism needs to be designed as large and costly devices such as high precision atomic clocks are not suitable for this scenario. Several different approaches have been utilized for T&S over long free space channels, including using pulsed lasers with various modulation patterns, and even combined with global positioning system (GPS) pulse per second (PPS) as a hybrid scheme. However, the ¿ 50 dB loss budget greatly increases the difficulty of precision beacon timing detection. A ultra-weak optical signal detection requires a detector with very high sensitivity and low noise level. A higher detecting bandwidth also provides a better timing jitter due to the sharper rise time. However, due to the limitations of the detector gain–bandwidth product (GBP), sensitivity and bandwidth need to be traded off to achieve the best SNR. Additionally, a large variation of incident power can result in pulse detection failure due to signal saturation. More importantly, the amplitude variation of the converted electronic pulse causes large discrimination errors of up to nanosecond level when using constant threshold signal discrimination, which significant increase the timing jitter and degrades the SNR [24]. Our solution is detailed below and in previous papers [15, 23].
@@ 2.1 Communication Link Design
The BB84 protocol encodes photons using four independent polarisation states, but the laser beam generated by an actual source can never reach 100% linear polarization. When photons coded as H-polarisation are incident on an ideal polarising beam splitter (PBS), a small fraction will inevitably enter into the V polarisation channel and thus contribute to a raw key error rate, denoted as $e_{SP}$. Similarly, the imperfection of practical PBS’s introduce unavoidable QBER as well denoted as $e_{PBS}$. In practice, $e_{PBS}$ can be estimated based on the extinction ratio of the PBS. The $e_{SP}$ is decided by the laser and polarizer. A polarizer with higher extinction ratio gives a lower QBER. These two parts are only determined by the device performance and working environment such as temperature and humidity, rather than the system design parameters, so they are defined as intrinsic QBER, $e_{I}$.
$e_{I}=F(e_{SP},e_{PBS})$ (1)
$F$ is the QBER combination function, which is defined in equation. 2. Assuming there are $n$ bits of binary variables (representing the polarisation state of photons) with random value. Part of the bits are flipped randomly in the first operation and another part of the bits are flipped randomly in a second operation. As long as the two operations sample randomly and independently, the actual flipped bits after two rounds can be calculated according to the inclusion-exclusion principle.
$F(e_{1},e_{2})=e_{1}+e_{2}-2e_{1}e_{2}$ (2)
@@
Where the $e_{A}$ is the part resulted from atmosphere, the $e_{N}$ is the part resulted from background noise and the $e_{DCR}$ is the noise comes from DCR. So, the whole QBER is defined as
$QBER=F(e_{I},e_{E})$ (4)
@@
In QKD security analysis, any QBER must be considered as resulting from the activity of an eavesdropper, which requires post-processing to perform privacy amplification of the key based on the QBER to ensure adequate security. Therefore, efforts to reduce the QBER can increase the key generation rate, and the optimization process of the system focuses on the $e_{E}$ since the $e_{I}$ is only determined by the performance of the off-the-shelf source, and the off-the-shelf discriminating optics in receiver (specifically referring PBS here) [31]. Bandpass filters are usually used to suppress background noise, which can greatly reduce all other noise beyond the band of the system signal. Generally speaking, the narrower filter window and the steeper cut-off edge, the better signal to noise ratio (SNR) could be achieved. However, considering the wavelength shift caused by the Doppler effect and temperature, the selection of the bandpass filter needs to be balanced between signal transmission and noise rejection to maximise the overall SKR. Besides, the background noise can be mitigated by gating or time-stamping the detection events. The minimum time interval for gating or time-stamping is set by the timing jitter of the detector, beacon and time-to-digital converter.
@@
As optical wavelengths are several orders of magnitude shorter than radio frequency signals, the diffraction limited divergence angles of 10 cm diameter beams are on the order of microradians. This brings optical communication the advantage of higher transmission efficiency compared to microwave communication, but it also leads to the need for an accurate pointing and tracking system to match the fields of view of satellites and OGS [32]. In addition, due to the relative motion of satellites with respect to ground station and the complex and changeable atmospheric environment, traditional clock synchronisation schemes for QKD in free space will lead to large jitter. The lack of synchronised high precision clocks results in inconsistencies between the transmitted and the received sequence and poor discrimination of background counts as a wide time gate is needed. Also, due to the continuous change of the satellite attitude relative to the OGS, providing an identical and stable polarisation reference is also essential to successful communication. Based on these considerations we have developed an elegant solution using a classical bright pulsed laser as beacon [15], and here is the first time it has be used for QKD demonstration. The pulses are designed to be highly synchronised (less than 15 ps) with the quantum signal to indicate the exact time of quantum signal arriving. Each two adjacent beacon pulses represent a bit and the HDBC is modulated onto the pulses train to provide a unique sequence index for each arriving pulse. Finally, the beacon source adopts linearly polarized light with a sufficiently high extinction ratio and is consistent with the polarization state of the H channel, so as to provide a reference for the OGS. Polarisation matching through satellite attitude control can solve the above problems, but it requires accurate, stable and real-time closed-loop feedback control, which dramatically increases the complexity of the system and increases the fuel consumption of the satellite. Additionally it cannot compensate for the error caused by any atmospheric effects. The most widely used method is to polarisation compensate the received signal with a set of automatically rotating waveplates. Using the linear polarisation properties of the beacon light, an algorithm will control the automatic rotation of two quarter-wave plates and one half-wave plate until the intensity of the received light reaches its maximum [^5] It is worth noting that because the beacon light and the signal light use different wavelengths, any birefringence caused by atmospheric effects is not independent of the wavelength, so it may lead to greater compensation errors in the signal at the receiving end..
3 System and Experimental Result